OpenStudy (anonymous):

Add. 1/2 + 3/4 + 1/2 A. 5/8 B. 5/4 C. 1 3/4 D. 1 7/4

OpenStudy (hennawho):

make a common denominator, which is 4 because 2 and 4 can go into it 1/2 = 2/4 3/4 = 3/4 1/2 = 2/4 2/4 + 3/4 + 2/4 add the numerators and keep the denominators the same 7/4 its the same thing as 1 3/4 because 4 x 1 + 3 = 7 and u put it over 4